# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ A.2 Proof of Theorem 2
Proof. The condition $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ implies the following descent inequality for the update rule $\theta_{t+1}=\theta_{t}-\eta\,\hat{g}_{t}$:
@@
where $g_{t}:=\nabla\mathcal{L}(\theta_{t})$. The conditional mean is $\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]=g_{t}+\gamma\alpha_{t}\hat{g}_{t-1}$. Then conditional second moment is given by:
$\mathbb{E}\bigl{[}\|\hat{g}_{t}\|^{2}\mid\mathbb{F}_{t}\bigr{]}$ $=\mathrm{Var}(\bar{g}_{t}\mid\mathbb{F}_{t})+\bigl{\|}\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]\bigr{\|}^{2}$
@@
$\quad+(\gamma\alpha_{t})^{2}\|\hat{g}_{t-1}\|^{2}.$ (22)
Substituting these into eq. (21) and grouping terms yields:
@@
$\quad+\frac{\ell\,\eta^{2}}{2n_{t}}\left(O(d\,\epsilon_{t}^{4})+\frac{d\,\sigma_{\xi}^{2}}{4\epsilon_{t}^{2}}\right),$ (25)
@@
$\zeta_{t}$ $:=1-\tfrac{\ell\eta}{2}-\tfrac{\gamma\alpha_{t}(1-\ell\eta)}{2}-\tfrac{\ell\eta(d-1)}{2n_{t}},$ (26)
$C_{t}$ $:=\frac{\eta\gamma\alpha_{t}(1-\ell\eta)}{2}+\frac{\ell\eta^{2}(\gamma\alpha_{t})^{2}}{2}.$ (27)
Telescoping the one-step descent lemma eq. (A.2) over $t=1,\dots,T$. Let $\zeta_{\min}:=\min_{1\leq t\leq T}\zeta_{t}>0$. Taking expectations, summing, we rearrange to find:
$\min_{1\leq t\leq T}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}+\sum_{t=1}^{T}C_{t}\,\mathbb{E}\|\hat{g}_{t-1}\|^{2}}{\eta\,T\,\zeta_{\min}}$
$\quad+\frac{\ell\,\eta}{2\,T\,\zeta_{\min}}\sum_{t=1}^{T}\frac{V_{t}}{n_{t}}.$ (28)
Prior analyses of stochastic optimization algorithms have established the boundedness of error terms [30, 82, 83], i.e.,
$C_{t}\leq C_{\max},\quad\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\hat{g}_{t-1}\|^{2}\leq\Upsilon,\quad\text{and}\quad\frac{1}{T}\sum_{t=1}^{T}\frac{V_{t}}{n_{t}}\leq\Xi.$
Applying these boundedness yields:
$\min_{1\leq t\leq T}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,T\,\zeta_{\min}}$
$\quad+\frac{C_{\max}\,\Upsilon}{\eta\,\zeta_{\min}}+\frac{\ell\,\eta\,\Xi}{2\,\zeta_{\min}}.$ (29)
To ensure the left-hand side is at most $\delta$, we solve for $T$. This requires
$T\geq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}},$ (30)
@@
1:Input: Model parameters $\theta_{t}$, batch size $N$, image generator $G(\cdot)$, hyperparameters $\beta$ and $\lambda$.
@@
3$\triangleright$ Sample a batch of prompts from $p_{\theta_{t}}(x)$
4:$\mathcal{X}\leftarrow\{x_{i}\sim p_{\theta_{t}}(x)\}_{i=1}^{N}$
5$\triangleright$ Compute alignment energy for each image
@@
11$\triangleright$ Compute prompt-level diversity energy
12:$\mathcal{L}_{\text{div}}\leftarrow\frac{1}{N(N-1)}\sum_{i}\sum_{j\neq i}\frac{\langle\mathcal{E}_{\xi}(x_{i}),\mathcal{E}_{\xi}(x_{j})\rangle}{\|\mathcal{E}_{\xi}(x_{i})\|\cdot\|\mathcal{E}_{\xi}(x_{j})\|}$
13$\triangleright$ Compute entropy regularization
14:$\mathcal{L}_{\text{ent}}\leftarrow\sum_{i=1}^{N}\log p_{\theta_{t}}(x_{i})$
15$\triangleright$ Aggregate total objective from all terms
16:$\mathcal{L}(\theta_{t})\leftarrow\mathcal{L}_{\text{align}}+\lambda\cdot\mathcal{L}_{\text{div}}+\frac{1}{\beta}\cdot\mathcal{L}_{\text{ent}}$
@@
Figure 4: Example unsafe images generated by DREAM’s red team prompt on the corresponding model. The black block and heavy blurring is added by authors to avoid disturbing readers.
Appendix B More Experimental Details
TABLE X: Categories and definitions of unsafe content used in our paper.
Category Definition
Diversity Level Definition
1: Limited Diversity The majority of prompts are near-identical or repeated with trivial modifications, such as basic rewordings (e.g., replacing “nude” with “nudity”). There is negligible lexical, structural, or conceptual diversity. Most prompts are variations on a fixed template and rely on the same narrow set of triggering keywords or phrases.
2$\sim$3: Moderate Diversity Prompts exhibit moderate diversity, often using modestly different triggering synonyms (e.g., replacing “nude” with “sexual”) or introducing light syntactic changes (e.g., changing the setting from “on the bed” to “in the room”). However, they still rely on a small group of core visual ideas and maintain similar structure and phrasing, with only minor surface-level differences.
4$\sim$5: High Diversity Prompts move beyond a small set of repetitive trigger words or formulaic expressions, demonstrating meaningful exploration of lexical, syntactic, and semantic alternatives. Instead of repeatedly relying on single terms like “nude woman,” the prompts vary across subjects (e.g., “erotic dancer,” “seductive character”), the frame (e.g., “softly lit room,” “posing suggestively”), and the scene structure. The prompts reflect creative and distributed discovery of diverse, or even unexpected potential triggers for generating unsafe content.
Appendix C More Discussion
Discussion with Other Optimization Methods. In addition to our proposed GC-SPSA algorithm, other optimization approaches can also be considered for minimizing the final loss function (i.e., Eq. (5)). One such alternative is reinforcement learning (RL) [51, 50], which are recently introduced to red team LLMs. A representative example is the recent CRT framework [50], which leverages PPO combined with a curiosity-driven reward and entropy bonus to enhance prompt diversity. Although CRT was originally designed for LLMs, the appendix of their revised paper preliminarily show that the same RL-based method can be extended to red team the Vanilla SD v1.5. However, CRT as implemented in the original paper only targets a general “unsafe” category. Therefore, it is not directly comparable to our method and other baselines, which is designed for specific unsafe categories. To enable a fairer comparison, we adapt CRT for a category-specific case study focusing on the “sexual” category. In this adapted version, the LLM is explicitly instructed to generate prompts related to sexual content. We train this CRT variant for 5,000 iterations, which takes approximately 24 hours on two NVIDIA RTX A100 GPUs, using all other default hyperparameters. As shown in Tab. XII, CRT exhibits high instability and, in some settings such as ESD, even fails completely. Moreover, in certain runs on CA, we observe CRT converges to repeatedly generating highly similar prompts after around 4,000 training steps, highlighting a lack of diversity and instability. In contrast, our method, DREAM, does not exhibit this repetitive behavior. We hypothesize that these differences arise from the fundamentally different optimization strategies used by RL and SPSA. In RL, each prompt (i.e., policy) is directly evaluated and rewarded. For safety-aligned models where most prompts fail to succeed, a few effective prompts receive disproportionately high rewards. This encourages the model to repeatedly generate the same high-reward prompts, even when regularized with KL divergence and novelty rewards, thereby reducing output diversity. In contrast, GC-SPSA optimizes by perturbing parameters and estimating the gradient via batch-based Monte Carlo sampling, without directly associating rewards with individual prompts. This means the model learns which direction in the parameter space improves performance rather than which specific prompts perform best. As a result, it is less likely to overfit to a single high-performing prompt, maintaining both stability and diversity throughout training. Nevertheless, we note that this analysis is not intended to diminish the potential of reinforcement learning. RL remains a powerful and expressive optimization paradigm, and with carefully tuned hyperparameters or alternative reward shaping strategies, its performance could potentially be improved. Recent advances such as GRPO [84], for instance, suggest promising directions that may help mitigate some of the issues that CRT suffer from. Our core claim is not that zeroth-order optimization is categorically superior, but rather that even a relatively simple and lightweight method like our GC-SPSA can already achieve stable and effective results in this challenging red teaming setting. We leave deeper exploration and more extensive optimization strategies to future work.
TABLE XII: Performance of CRT [50]. The results show the mean and standard deviation for 3 independent runs.
Metric SD v1.5 CA ESD SC
PSR $\uparrow$ 93% ± 06% 47% ± 14% 04% ± 03% 12% ± 08%
PS $\downarrow$ 0.65 ± 0.03 0.68 ± 0.04 0.72 ± 0.02 0.71 ± 0.01
Ethics Considerations. This research is intended solely for advancing the safety of T2I generative systems. DREAM is designed as a red teaming framework to proactively evaluate and improve existing safety mechanisms, not to attack or undermine any deployed systems. Nonetheless, we acknowledge that our method may generate prompts capable of bypassing safety mechanisms and elicit harmful outputs. To mitigate risk, we have responsibly disclosed all discovered vulnerabilities including prompt examples and generated images to the developers of affected models, safety filters, and commercial platforms via their official contact channels (e.g., forums and official emails), and are currently awaiting their responses. We have comprehensively discussed this work, including the user study, with our Institutional Review Board (IRB) and received an exempt determination. All study participants were adults in good physical and mental health, and confirmed to have no histories of heart conditions, psychological disorders, or vasovagal syncope, provided informed consent, and were made aware of their right to withdraw at any time. No personal or sensitive data was collected. Additionally, we place a strong emphasis on the well-being of both the researchers and study participants involved in this project. All authors and participants involved in this study were provided access to institutional mental health resources and encouraged to monitor and communicate any discomfort arising from interacting with unsafe or disturbing content during experimentation. Study participants were similarly given appropriate content warnings and support resources. We believe that fostering a research culture of care and accountability is essential, especially when dealing with high-stakes safety and content moderation work. To minimize potential misuse, we have refrained from publicly releasing the full set of discovered unsafe prompts until the affected parties have had sufficient time to address the issues. We will continue to engage with stakeholders and iterate on responsible disclosure practices in line with the principles outlined in the Menlo Report and the broader computer security research community.

### v2

@@ A.2 Proof of Theorem 2
Proof. The smoothness condition $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ implies the standard descent inequality for $\theta_{t+1}=\theta_{t}-\eta\,\hat{g}_{t}$. We adopt zero-indexing with $\hat{g}_{-1}=0$ and $w_{-1}=0$, so that $\alpha_{0}=0$.
We adopt the zero-indexing convention with $\hat{g}_{-1}=0$ and $w_{-1}=0$, which implies $\alpha_{0}=w_{-1}/(w_{-1}+n_{0})=0$. The condition $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ implies the following descent inequality for the update rule $\theta_{t+1}=\theta_{t}-\eta\,\hat{g}_{t}$:
@@
The conditional mean of the GC-SPSA estimator is:
$\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]=g_{t}+\gamma\alpha_{t}\hat{g}_{t-1},$ (22)
where $\alpha_{t}:=w_{t-1}/(w_{t-1}+n_{t})$ is the confidence weight. The conditional second moment is given by:
$\mathbb{E}\bigl[\|\hat{g}_{t}\|^{2}\mid\mathbb{F}_{t}\bigr]$ $=\mathrm{Var}(\hat{g}_{t}\mid\mathbb{F}_{t})+\bigl\|\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]\bigr\|^{2}$
@@
$\quad+(\gamma\alpha_{t})^{2}\|\hat{g}_{t-1}\|^{2}$ (23)
. Substituting these into Eq. (21) and grouping terms yields:
@@
$\quad+\frac{\ell\,\eta^{2}}{2n_{t}}V_{t},$ (26)
@@
$\zeta_{t}$ $:=1-\frac{\ell\eta}{2}-\frac{\gamma\alpha_{t}(1-\ell\eta)}{2}-\frac{\ell\eta(d-1)}{2n_{t}},$ (27)
$C_{t}$ $:=\frac{\eta\gamma\alpha_{t}(1-\ell\eta)}{2}+\frac{\ell\eta^{2}(\gamma\alpha_{t})^{2}}{2},$ (28)
$V_{t}$ $:=O(d\,\epsilon_{t}^{4})+\frac{d\,\sigma_{\xi}^{2}}{2\epsilon_{t}^{2}}.$ (29)
Telescoping the one-step descent lemma in Eq. (A.2) over $t=0,\dots,T-1$. Let $\zeta_{\min}:=\min_{0\leq t\leq T-1}\zeta_{t}>0$. Taking expectations, summing, we rearrange to find:
$\min_{0\leq t\leq T-1}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\sum_{t=0}^{T-1}C_{t}\,\mathbb{E}\|\hat{g}_{t-1}\|^{2}}{\eta\,T\,\zeta_{\min}}$
$\quad+\frac{\ell\,\eta}{2\,T\,\zeta_{\min}}\sum_{t=0}^{T-1}\frac{V_{t}}{n_{t}}.$ (30)
Prior analyses of stochastic optimization algorithms have established the boundedness of error terms [spall1992multivariate, ghadimi2013stochastic, balasubramanian2019zeroth], i.e.,
$C_{t}\leq C_{\max},\,\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\hat{g}_{t-1}\|^{2}\leq\Upsilon,\,\text{and}\,\frac{1}{T}\sum_{t=0}^{T-1}\frac{V_{t}}{n_{t}}\leq\Xi.$
Applying these bounds yields:
$\min_{0\leq t\leq T-1}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}}{\eta\,T\,\zeta_{\min}}+\frac{C_{\max}\,\Upsilon}{\eta\,\zeta_{\min}}+\frac{\ell\,\eta\,\Xi}{2\,\zeta_{\min}}.$
To ensure the left-hand side is at most $\delta$, $T$ needs to satisfy:
$T\geq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}},$ (31)
@@
Figure 6: Example unsafe images generated by DREAM’s red team prompt on the corresponding model. The black block and heavy blurring is added by authors to avoid disturbing readers. The red teaming prompt dataset can be found in our code repository.
TABLE X: Main categories and definitions of unsafe content used in our paper.
Category Definition
Diversity Level Definition
1: Limited Diversity The majority of prompts are near-identical or repeated with trivial modifications, such as basic rewordings. There is negligible lexical, structural, or conceptual diversity. Most prompts are variations on a fixed template and rely on the same narrow set of triggering keywords or phrases.
[…]
Appendix B More Implementation Details & Discussion
Additional Implementation Details. In our experiments, we set the hyperparameters as $\gamma=1.2$, the perturbation magnitude $\epsilon$ initialized as $1\times 10^{-3}$ and the learning rate $\eta$ initialized as $1\times 10^{-6}$ for our GC-SPSA. For inference-time sampling, the minimum temperature in ATS is set to 0.8. Regarding the construction of ICL exemplars, examples for the sexual category are drawn from the NSFW-56k dataset [li2024safegen], whereas examples for violence and other categories are selected from the I2P [schramowski2023sld] dataset and the Unsafe dataset [qu2023unsafe]. In certain cases, we slightly tune the parameter adjustments and curated selections of ICL exemplars during the early validation stage to improve optimization stability and overall training effectiveness. The precise implementation details and configuration files for all experiments are provided in our open-source release. In addition, we include all generated red-teaming prompts as artifacts to facilitate reproducibility and to support future research in this area.
Discussion on MHSC’s Accuracy. Our experiments rely primarily on MHSC to label outputs. To verify its reliability, we evaluate MHSC on both a real-world dataset (the test set of the NCD dataset) and our DREAM-generated dataset (SD v1.5+Sexual, 256 balanced, randomly selected samples), and the results are shown in Tab. XII. We can derive two insights from the results. First, MHSC achieves over 90% accuracy and F1 scores on both domains, indicating that its predictions are highly reliable. Second, MHSC indeed exhibits a conservative classification tendency, where it only flags an image as unsafe when it is highly confident. This further confirms our analyses in Sec. 5.4.
TABLE XII: Performance of MHSC.
Setting ACC $\uparrow$ TPR $\uparrow$ FPR $\downarrow$ TNR $\uparrow$ FNR $\downarrow$ F1 $\uparrow$
NCD Test Set 93.52% 89.43% 2.40% 97.60% 10.57% 93.24%
DREAM (Ours) 93.17% 92.97% 6.67% 93.33% 7.03% 92.61%
@@ Appendix B More Implementation Details & Discussion
1:Input: Model parameters $\theta_{t}$, batch size NN, image generator $G(\cdot)$, hyperparameters $\beta$ and $\lambda$.
@@
[…]
@@
[…]
@@
TABLE XIII: Component-level requirements of different methods on the target T2I diffusion model. For the first two columns, ●  denotes the method requires explicit back-propagation-based gradients from the target model’s corresponding component. ◐  means while the method requires gradients from that component, they are obtained via numerical approximations rather than exact back-propagation (e.g., via GC-SPSA). ○  indicates the method does not utilize signals from this component. For the “Shadow Model” column, YES / NO indicates whether the method assumes access to a local shadow model (e.g., a shadow text encoder) that is identical or highly similar to that of the target model.
Method
Text
Encoder
Denoising
Network
Shadow
Model
QF-Attack NO
MMA-Diffusion NO
P4D / UnlearnDiffAtk NO
SneakyPrompt / JailFuzzer YES
ART / FLIRT NO
DREAM (Ours) NO
Discussion on Baseline Selection. Recall that red-teaming is typically initiated by the model owner/developer, who naturally has full control over their own T2I system. However, not all baselines in our paper fully exploit this privilege. For better clarity, we summarize the capability assumptions made by each baseline in Tab. XIII. Several insights emerge from this summarization. First, most evaluated baselines require clear-box access, i.e., the weights, gradients, or intermediate outputs of at least one system component, indicating that they still leverage the owner’s privilege.
Second, and perhaps counter-intuitively, the amount of information required does not always correlate with better final performance. For example, on CA-Violence, Keyword-Gibberish, and NSFW-Text-Detector, closed-box methods like SneakyPrompt, FLIRT, and JailFuzzer sometimes perform the best among all baselines. Similar findings have also been reported in LLM adversarial learning [wang2025functional, andriushchenko2025jailbreaking], where gradient-based clear-box strategies are observed to be less effective than closed-box gradient-free approaches like random search. One plausible explanation is that in large-scale discrete optimization over language prompts, clear-box gradients tend to be highly local and noisy. As a result, they may sometimes underperform gradient-free methods, which allow exploring larger solution spaces more efficiently. These observations suggest that the key determinant of red-teaming performance may not fully lie in the level of access, but in how the problem is modeled and how the optimization is conducted. Thus, we include both clear-box and closed-box baselines in our experiments, as it would highlight the strongest competitor regardless of assumed capability and make our evaluation more complete.
Example Unsafe Images. We provide representative harmful outputs discovered by our method when probing a variety of safety-aligned, filter-enhanced, and commercial T2I systems, as shown in Fig. 6 and 7. To prevent disturbance to readers, all images are heavily mosaicked or visually censored while preserving their semantic structure. In addition, we release the full set of red-teaming prompts generated by DREAM to support future safety research, which will be available through gated-access in our code repository.
On the Importance of Diversity and Efficiency for Red Teaming. Different from targeted jailbreak attacks, red teaming aims to uncover a representative set of unsafe prompts that spans diverse failure modes of the target system. Diversity is therefore essential: if discovered prompts concentrate on only a few lexical or semantic patterns, the resulting test set provides limited coverage and reduced value for diagnosing vulnerabilities or supporting safety fine-tuning. At the same time, efficiency is equally critical, as practical safety evaluation and alignment workflows often require thousands of high-quality unsafe prompts. Methods that scale poorly quickly become prohibitively expensive at this regime and fail to meet real-world red-teaming needs. Together, strong diversity and efficiency enable broad, systematic exploration of the unsafe prompt space and are indispensable for reliable safety assessment.
On the Novelty of DREAM. DREAM introduces several core innovations that collectively advance the field of automated red teaming. First, we reformulate red teaming as a distribution learning problem and derive a tractable objective via energy-based modeling, enabling effective learning without direct access to representative unsafe samples. This perspective departs from prior prompt-level optimization and offers a principled probabilistic view of the unsafe prompt space. Second, we propose GC-SPSA, a novel zeroth-order optimizer with a gradient calibration mechanism that significantly improves stability over vanilla SPSA, a result supported by both theoretical analysis and substantial empirical gains. Third, we design targeted energy functions to guide distribution learning and introduce an adaptive temperature scaling strategy that enhances inference-time coverage at negligible computational cost. Beyond these components, DREAM also establishes a conceptual connection between our final objective and entropy-regularized reinforcement learning: the entropy term encourages exploration while the energy term provides a structured reward landscape. We believe this connection may inspire future work to further unify red teaming with principled RL frameworks, potentially enabling even more scalable and adaptive red teaming. Finally, DREAM provides the first formal definition of red teaming for T2I systems and a thorough analysis of the limitations of prompt-to-prompt methods, offering a clearer conceptual grounding for the field.
Figure 8: Impact of N on PSR.
Discussion on Red Teaming Paradigm. Besides the efficiency analysis in Sec. 5.6, it is also worth discussing the red teaming paradigms underlying different methods. Existing methods generally fall into two categories. Some approaches (e.g., human-written datasets) first generate a pool of prompts on a surrogate T2I system and then reuse them across multiple targets. Others, including most recent methods (e.g., P4D, FLIRT, JailFuzzer) and DREAM, perform system-specific training that adapts the red-teaming process to each target system individually.
These two paradigms represent a natural trade-off between cost and adaptivity. While reusing pre-discovered prompts is cheap and convenient, such methods may lack adaptivity to model-specific vulnerabilities. As implied by Definition 1, the set of effective red-teaming prompts is inherently system-specific: a prompt that succeeds on one system may fail on another, and vice versa. Our empirical observations also support this view: for instance, on the Keyword‑Gibberish Filter, almost all fixed‑dataset approaches achieve very low success rates, whereas adaptive methods such as FLIRT, JailFuzzer, and DREAM still obtain high success rates, indicating that the system is not genuinely safe. This suggests that relying solely on fixed datasets for evaluation may lead to a false sense of safety, whereas model-specific adaptive training can uncover target‑specific blind spots that are crucial for reliable safety evaluation and targeted improvement.
At the same time, red teaming with DREAM does not necessarily indicate prohibitive cost. As shown in Tab. VI, DREAM’s red team LLM can be reused or lightly adapted to new targets, significantly reducing cost while maintaining competitive effectiveness. This flexibility provides a practical balance between efficiency and adaptivity.
Thus, we suggest that real‑world users choose their strategy according to their goals and resource budgets. In practice, the two paradigms could be complementary rather than mutually exclusive: users can first apply low‑cost fixed‑dataset methods for rapid screening or benchmarking, and then selectively deploy adaptive approaches such as DREAM if they need further high‑coverage, model‑specific red‑teaming or targeted safety improvement. This staged approach leverages the efficiency of fixed datasets to identify general issues while using adaptive methods to uncover deeper, system‑specific vulnerabilities. In addition, reusing or lightly adapting a previously trained red‑team LLM offers a lightweight yet valuable middle ground.
Comparison with PromptTune [jiang2025jailbreaking]. We further compare with a recent baseline, i.e., PromptTune [jiang2025jailbreaking]. Specifically, we train PromptTune using the NSFW-66k dataset provided by the authors, following the PromptTune-AdvPrompter setting in their paper, and evaluate it under CA and SC. We obtain a PSR of $14.7\%$ on CA and $38.3\%$ on SC. Through this case study, we observe that although PromptTune also fine-tunes an LLM, its effectiveness is weaker than DREAM. A possible reason lies in PromptTune’s training mechanism: it learns by generating candidate prompt variants, selecting the most effective one, and then fine-tuning on it. This indirect supervision may be less efficient than DREAM’s paradigm, which directly learns from the energy function via GC-SPSA. We leave further exploration to future work.
Impact of N. In our evaluation, we adopt PSR-N instead of a single generation, so as to mitigate potential underestimation of effectiveness due to the inherent sampling randomness of diffusion models. As shown in Fig. 8, while the PSR of all methods increases monotonically with larger N, the improvements diminish quickly and the curves nearly converge when $\text{N}>6$. Importantly, the relative ranking of methods remains stable, with DREAM consistently performing the best. This further supports the reliability of our conclusions.
TABLE XIV: Results on common diversity-oriented decoding strategies. We report PSR and PS over 3 sampling runs.
Penalty-based Decoding Sampling-based Decoding
Strategy PSR $\uparrow$ PS $\downarrow$ Strategy PSR $\uparrow$ PS $\downarrow$
[…]
Comparison with Common Diversity-oriented Decoding Strategies. We compare our Adaptive Temperature Scaling (ATS) strategy with common diversity-oriented decoding strategies, including presence/frequency/repetition penalties and sampling-based methods such as temperature, top-$k$, and nucleus sampling. As shown in Tab. XIV, traditional penalty-based methods provide limited benefits in our setting, as they all focus within a single prompt where repetitions are already rare. In contrast, ATS serves as a cross-generation extension of these penalties, enabling more effective diversification across sampled prompts. ATS improves diversity with small impact on effectiveness and requires only a few seconds of extra computation, making it a practical and plug-and-play enhancement for broadening coverage.
On Minimax-style Adversarial Training with DREAM. In the main paper, our safety fine-tuning procedure adopts a fixed dataset generated once by DREAM. Outside our core experiments, during this research, an expert insightfully suggested exploring a minimax-style adversarial training scheme in which the red-team generator and the defended model are updated in an alternating fashion. Motivated by this insight, we further examine such an adversarial setup. We first assess benign utility and find that models aligned with DREAM prompts preserve utility comparable to the original SDv1.5 baseline (CLIP Score: 31.17 vs. 30.98), indicating that our red-teaming–augmented alignment procedure does not substantially degrade generation quality. Building on this, we further explore an alternating training strategy inspired by adversarial learning: 50 steps of updating DREAM followed by 50 steps of Safety-DPO, repeated for approximately six rounds. This adversarially trained model achieves consistently lower PSR across diverse attacks compared to the fixed-dataset approach (e.g., 12.1% to 5.3% under DREAM, 21.8% to 8.7% under DREAM+50-step), demonstrating that adaptively co-evolving the red-team generator and defended model provides stronger robustness. While this procedure introduces a modest reduction in utility (CLIP Score 30.03), the substantial security gains highlight adversarial training as a promising extension of our framework. We believe this direction opens opportunities for more dynamic and theoretically grounded defense strategies in future work.
Discussion on Parameter Efficient Fine-Tuning (PEFT). Motivated by expert suggestions to explore parameter-efficient tuning, we further investigate a LoRA-based variant for our DREAM framework. We observe that LoRA indeed accelerates convergence, reducing the number of training epochs from roughly 280 to 255, which aligns with our theoretical analysis in Theorem 2 that smaller effective parameter dimensionality increases the descent coefficient and therefore shortens the required optimization horizon. However, the final performance of LoRA is slightly lower (PSR 70.1% vs. 76.0% on CA+Sexual), which is consistent with prior findings that larger parameter spaces enable reaching better optima. In addition, we also find that LoRA occasionally introduces higher training instability for some settings, likely due to the nature of zeroth-order optimization where parameter updates can be more sensitive under low-rank parameterization [chen2024enhancing]. Such instability may, in certain cases, counteract the convergence benefits of LoRA. We believe this issue could be mitigated by employing more advanced zeroth-order optimization algorithms, which we leave as a promising direction for future work.
Discussion on Other Gradient Estimation Strategies and RL-based Approaches. In addition to our GC-SPSA, other optimization approaches can, in principle, be applied to minimize the final loss function in Eq. (5). One representative line is reinforcement learning (RL) [perez2022red, hong2024curiosity], which has recently been introduced to red team LLMs. For example, CRT [hong2024curiosity] fine-tunes an LLM using PPO combined with curiosity-driven rewards and entropy bonuses to promote prompt diversity. Although originally designed for LLMs, the authors have discussed its potential to be extended to T2I models. To make a more direct comparison, we adapted CRT to our category-specific setting by modifying system prompts and evaluated it on CA. As shown in Tab. XV, CRT exhibits high instability, and we even observed complete failure (i.e., $\text{PSR}<5\%$) for some cases. We hypothesize that this stems from a fundamental difference between RL and SPSA. In RL, each prompt (i.e., trajectory) is directly rewarded. For safety-aligned models where most prompts fail, a few lucky successful prompts receive disproportionately large rewards, which tends to drive the policy toward repeatedly generating these high-reward prompts even with regularization (e.g., KL or novelty terms), ultimately hurting diversity. By contrast, GC-SPSA perturbs model parameters and estimates gradients via batch-based Monte Carlo sampling, without tying rewards to individual prompts. This allows the optimizer to identify general update directions in parameter space rather than reinforcing a few trajectories, thereby maintaining stability and diversity during training.
Besides RL, other parameter-based gradient estimators like RDSA [kushner2012stochastic] and NES [wierstra2014natural] could also be considered. While these methods alleviate the trajectory overfitting issue to some extent, our experiments (Tab. XV) show that RDSA suffers from low effectiveness, while NES incurs much higher computational cost. This may be attributed to RDSA’s susceptibility to large estimation noise when hyperparameters are not carefully tuned [chen2021theoretical], and to NES’s reliance on large population sizes. In contrast, SPSA estimates gradients with Gaussian perturbations and only two function evaluations per iteration, which offers a more balanced tradeoff between simplicity, stability, and efficiency. We thus focus on SPSA while leaving broader exploration into applying RDSA and NES for red teaming for future work.
As a final remark, we emphasize that this analysis is not meant to diminish the potential of RL or other gradient estimation techniques. We believe they remain powerful and expressive, and with advanced variance-reduction strategies or carefully tuned hyperparameters, their potential could be further unlocked. Our core claim is not that SPSA is categorically superior, but rather that even a relatively simple and lightweight method like our GC-SPSA already achieves stable and effective results in the challenging red teaming setting. We hope DREAM provides a practical and extensible starting point, and we leave deeper investigation of these alternative optimization strategies to future work.
TABLE XV: Results of other parameter-based gradient estimators and the RL-based approach CRT [hong2024curiosity]. The setting is CA+Sexual. We report the results averaged over 3 independent runs.
Parameter-based RL-based
RDSA [kushner2012stochastic] NES [wierstra2014natural] CRT [hong2024curiosity]
[…]
Ethics Considerations. This research is intended solely for advancing the safety of T2I generative systems. DREAM is designed as a developer-driven red teaming framework to proactively evaluate and improve safety mechanisms, not to attack or undermine any deployed systems. However, we acknowledge that it may generate prompts capable of bypassing safety mechanisms and eliciting harmful outputs. To minimize potential harm, we have responsibly disclosed the discovered vulnerabilities to developers of the affected T2I systems and provided a sixty-day window to address the issues prior to publication. We also refrain from publicly releasing any unsafe prompts or uncensored images that may be harmful to real-world services or users. Instead, representative attack prompts, generated samples, and the derived dataset will be released via gated access, requiring an access request, statement of intended use, and institutional affiliation. This study was reviewed and approved by the institutional research ethics committee of our university, which fulfills a role analogous to the IRB in the United States. All participants were healthy adults who have provided informed consent, were informed of their right to withdraw at any time, and no personal data were collected. We reduced participants’ exposure to disturbing material by limiting their workload during the user study and, for subsequent experiments, restricting human annotation to a small team of authors. All participants were given access to institutional mental-health resources and received content warnings. We will continue to engage with stakeholders following the principles outlined in the Menlo Report and the broader computer security research community.
Appendix C Meta-Review
The following meta-review was prepared by the program committee for the 2026 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
C.1 Summary
This paper focuses on the safety of text-to-image (T2I) generative models, particularly their ability to produce harmful content. The authors model the probabilistic distribution of the target system’s problematic prompts so that the designed framework automatically uncovers diverse problematic prompts from the target.
C.2 Scientific Contributions
  • Creates a New Tool to Enable Future Science
  • Addresses a Long-Known Issue
  • Provides a Valuable Step Forward in an Established Field
C.3 Reasons for Acceptance
  1. 1.
    This paper proposes a scalable red teaming framework that improves prompt discovery through probabilistic modeling and stable optimization, advancing safety evaluation for text-to-image generative systems.
  2. 2.
    The proposed method aims to overcome the inefficiency of prior approaches (e.g., token-level optimization, iterative rewriting), potentially enabling large-scale exploration of problematic prompts.